Razor:

Can someone tell me if I'm right? I have 10 pictures, I will post one for each one answered thankz o-o mostly calling on dude

Razor:

1 attachment
dude:

Looks right

Razor:

1 attachment
Razor:

eeeeeeeeeee

dude:

Wait why did you cross out A

dude:

You flipped the numbers \(\large \frac{9-6}{-5-(-3)}=\frac{3}{-5+3}\)

Razor:

A = M = -3/2

Razor:

That's option A...

dude:

Yeah, that's why I asked your decision to cross it out

Razor:

Because I was doing the problem with my mom and she said it was wrong o-o

dude:

Maybe she mixed up le numbers \((-3,6)=(x_1,y_1)\) \((-5,9)=(x_2,y_2)\) Use the equation: \(\huge \frac{y_2-y_1}{x_2-x_1}\)

dude:

Which would give you this (I did this above) \(\large \frac{9-6}{-5-(-3)}\)

Razor:

whats the 1 by the x at the top for?

dude:

The subscript (The numbers below a letter) just represent which point it is So \(x_1\) just means the first x value \(x_2\) means the second x value This is meant to keep the order consistent, if you subtract the wrong way, you'll get a completely different answer You always want to subtract the second point from the first one

Razor:

Oh, okay.

dude:

yeah

Razor:

1 attachment
Razor:

Just solves those?

dude:

Yes

Razor:

like 9-6 and then / then -5 - (-3) and then make that a fraction

dude:

Yes

Razor:

Okie, brb.

Razor:

oof, su 3/-2

dude:

Yes Since there is only a negative in the denominator the whole equation is negative \(-\large \frac{3}{2}\)

Razor:

okay. so it was option A this whole time XD Glad I double checked

Razor:

1 attachment
dude:

Not quite So you added 4 and 2 But you're supposed to add 9 and 7 as well (which is not 17)

Razor:

16 XD

dude:

Yes \(\large \frac{6}{16}\) They can both be divided by something, do you know what that is?

Razor:

3/7 lol

dude:

If that were true it would be in the answer choices \(\large\frac{3}{8}\)* xD

Razor:

oof XD

Razor:

I waz looking at the wrong clue LOL okie 3/8... check

Razor:

1 attachment
dude:

It's rise/run It goes up 1, not 2

Razor:

flip e,e? ಠ_ಠ

dude:

Yes

Razor:

Su glad I'm double checking these,

Razor:

1 attachment
dude:

Looks right

Razor:

1 attachment
Razor:

kayz just fly through them lol

dude:

Right

Razor:

1 attachment
dude:

You had already posted this?

Razor:

oof oopsie

dude:

No problem

Razor:

1 attachment
Razor:

Dun't know if yew can read it or not

dude:

I can read it, the slope is just the number in front of the x Nothing to alter there, so it's not A

Razor:

alter?

dude:

Nothing to change* It's just what you see in front of the x in the equation \(\large y=\color{red}{\frac{1}{8}}x=1\)

Razor:

ohhh... so its just gonna be 1/8 then? XD

dude:

Yes

Razor:

Yew should honestly become a math teacher e,e

dude:

Haha, I suck at teaching.

Razor:

1 attachment
Razor:

Doing well su farz XD

dude:

Right

Razor:

1 attachment
dude:

Right

Razor:

eeeeeeeeeeeeeeeeeeee

dude:

Woot woot

Razor:

1 attachment
dude:

Correct

Razor:

Thats all ten o-o

Razor:

Kay thankz o--o

dude:

Good job boi No problemo